# Distributionally robust monopoly pricing: Switching from low to high prices in volatile markets

### v1

@@ 6 Low and high pricing in service operations management
We next consider monopoly pricing for delay-prone services, a setting that is technically more involved than the classical monopoly pricing problem considered so far, but again leads to switching from low to high prices when the risk becomes large enough.
Consider a monopolistic service firm that operates as an $M/M/1$ queue. Potential customers arrive according to a Poisson process with arrival rate $\lambda$. Customers who decide to join the queue are served in order of arrival by a single server. Service times are exponentially distributed with rate $\theta$. Customers have individual service valuations that are modeled as i.i.d. random variables. In traditional models the valuation distribution is assumed to be known to the system manager, an assumption that we will relax below. Customers know their valuation but cannot observe the queue length, and decide whether to join or not only based on their assessment of the expected waiting times. Such queues are called unobservable queues and widely studied in the field of rational queueing theory; see Hassin and Haviv (2003) for an extensive overview.
For notational consistency, denote the random valuation of service by $X$, so that a customer only joins the queue when this $X$ exceeds the sum of price $p$ and holding costs $hW(\gamma)$ with $W(\gamma)$ some continuous function that increases with $\gamma$, where a standard choice would be $W(\gamma)=\frac{\gamma}{\theta(\theta-\gamma)}$, the expected waiting time in an $M/M/1$ queue with effective arrival rate $\gamma$ and service rate $\theta$. The parameter $h$ is the cost of waiting per time unit and measures the customer’s delay sensitivity. Let $\mathbb{P}$ denote the distribution of $X$. An arbitrary customer then joins with probability $\mathbb{P}(X>p+hW(\gamma))$. (For mathematical convenience we use a strict inequality $>$ instead of $\geq$. Note that by Lemma 3.1 this will not affect the worst-case tail bound for the ambiguity sets considered in this paper.) Since customers arrive according to a Poisson process with rate $\lambda$, the effective arrival rate $\gamma=\gamma(p,\mathbb{P})$ solves
@@
A revenue-maximizing seller seeks to solve
$\sup_{p}p\cdot\gamma(p,\mathbb{P})$ (40)
and we consider the robust setting where the seller does not know the distribution $\mathbb{P}$ of $X$, but instead only knows that the valuation distribution lies in the ambiguity set $\mathcal{P}$. Taking a pessimistic approach, the seller then sets a deterministic price that maximizes the expected revenue under the worst-case scenario.
It is not guaranteed that (39) has a solution for given $p$ and $\mathbb{P}\in\mathcal{P}$. We therefore set $\gamma(p,\mathbb{P})=\infty$ if no solution to (39) exists. We assume that for every $p$, there is at least one $\mathbb{P}\in\mathcal{P}$ for which $\gamma(p,\mathbb{P})<\infty$. This is in particular true for ambiguity sets satisfying the two conditions in Theorem 43, as becomes clear from the proof. We define
@@
and then the objective of a robust seller is to solve
$\sup_{p}\,p\cdot\gamma^{*}(p,\mathcal{P}).$ (42)
Theorem 6.1.
@@
    $\inf_{\mathbb{P}\in\mathcal{P}}\mathbb{P}(X>p+hW(x))$ is attained by some distribution $\mathbb{P}_{x}$ for every $x$.
@@
Proof. Recall that $W$ is an increasing function and so if $x<x^{\prime}$, then also $p+hW(x)<p+hW(x^{\prime})$. This implies that $\lambda\cdot f(x)$ is non-increasing in $x$. Because $g(x)=x$ is an increasing function in $x$, it follows that the equation
$x=\lambda\cdot\inf_{\mathbb{P}\in\mathcal{P}}\mathbb{P}(X>p+hW(x))$ (44)
has a unique solution $x^{*}$. We will next argue that $x^{*}\geq\gamma^{*}$. Recall that $\gamma^{*}=\inf_{\mathbb{P}\in\mathbb{P}}\gamma(p,\mathbb{P})$ with $\gamma(p,\mathbb{P})$ the unique solution to the equation
$\gamma=\lambda\cdot\mathbb{P}(X>p+hW(\gamma))$ (45)
for the given distribution $\mathbb{P}$. Because of the second assumption we know that there is a distribution $\mathbb{P}_{x^{*}}$ so that
$x^{*}=\lambda\cdot\inf_{\mathbb{P}\in\mathcal{P}}\mathbb{P}(X>p+hW(x^{*}))=\lambda\cdot\mathbb{P}_{x^{*}}(X>p+hW(x^{*})).$ (46)
This means that $\gamma^{*}\leq x^{*}<\infty$.
We will next argue that $\gamma^{*}=x^{*}$ using a proof by contradiction. Suppose that equality does not hold, i.e., $x^{*}>\gamma^{*}$. Then there exists a distribution $\mathbb{P}^{\prime}\in\mathcal{P}$ where $\gamma^{\prime}=\gamma^{\prime}(p,\mathbb{P}^{\prime})$ satisfies
$\gamma^{\prime}=\lambda\cdot\mathbb{P}^{\prime}(X>p+hW(\gamma^{\prime}))$ (47)
and $\gamma^{*}\leq\gamma^{\prime}<x^{*}$. Because the function $x\mapsto\mathbb{P}^{\prime}(X>p+hW(x))$ is non-increasing, it follows that
$\mathbb{P}^{\prime}(X>p+hW(x^{*}))\leq\mathbb{P}^{\prime}(X>p+hW(\gamma^{\prime}))=\frac{\gamma^{\prime}}{\lambda}<\frac{x^{*}}{\lambda}=\mathbb{P}_{x^{*}}(X>p+hW(x^{*})).$ (48)
By definition of $\mathbb{P}_{x^{*}}$, i.e., the second assumption of the theorem statement, this then implies that
$\mathbb{P}^{\prime}(X>p+hW(x^{*}))<\inf_{\mathbb{P}\in\mathcal{P}}\mathbb{P}(X>p+hW(x^{*})),$ (49)
which is a contradiction by definition of the infimum. This shows that $x^{*}=\gamma^{*}$ and completes the proof. ∎
@@
Lemma 3.1 shows that Theorem 43 applies to all ambiguity sets considered in this paper: The function (6) is always continuous in $p$, and worst-case distributions are identified in the proof of Lemma 3.1 in Appendix A. Take as example the ambiguity set $\mathcal{P}(\mu,\beta)$ of all distributions with given mean $\mu$ and maximal valuation $\beta$. Lemma 3.1 with $\underaccent{\bar}{\sigma}=0$ and $\bar{\sigma}=\sqrt{\mu(\beta-\mu)}$ then yields
$\inf_{\mathbb{P}\in\mathcal{P}(\mu,\beta)}\mathbb{P}(X>x)=\frac{\mu-x}{\beta-x},\quad x\in(0,\mu),$ (50)
and hence, by taking $x=p+hW(\gamma^{*})$, we have that $\gamma^{*}$ solves
$\frac{\mu-p-hW(\gamma^{*})}{\beta-p-hW(\gamma^{*})}=\frac{\gamma^{*}}{\lambda}.$ (51)
For $W(\gamma)=\frac{\gamma}{\theta(\theta-\gamma)}$ the worst-case arrival rate $\gamma^{*}$ can be obtained in closed-form. Moreover, for this case, we can also derive an explicit upper bound on $\gamma^{*}$ using the fact that $\mu-p-hW(\gamma^{*})$ needs to be nonnegative. This implies that
$\gamma^{*}\leq\gamma_{\max}\mathrel{\mathop{:}}=\theta-\frac{\theta h}{\theta\mu+h-\theta p},$ (52)
and this bound becomes exact in the large market size limit $\lambda\to\infty$. This also gives an upper bound $p\gamma_{\max}$ on the worst-case revenue, which is maximized by
$p_{\max}\mathrel{\mathop{:}}=\mu-\frac{\sqrt{h^{2}+\theta h\mu}-h}{\theta}.$ (53)
@@
Pricing in this service system depends on the delay function $W(\cdot)$, the service rate $\theta$ and the additional parameters $h$ and $\theta$. Observe that for $h=0$ we retrieve the classical monopoly pricing problem. See Figure LABEL:plotAB for an example. Knowing the mean and maximal value thus leads to a unique optimal price that decreases with $\theta$. As in the original monopoly pricing problem with information on the mean and maximal value, there is no low-high pricing phenomenon. The result, however, is interesting in its own right, as distribution-free pricing for these service systems has not been covered in the literature.
Now assume the seller knows the mean $\mu$, bound on support $\beta$ and the standard deviation $\sigma$. Theorem 43 implies that the worst-case tail bound from Lemma 3.1 for $\underaccent{\bar}{\sigma}=\bar{\sigma}$ can be used to solve the max-min pricing (42). Figure LABEL:FigmainhQ3 shows how the arrival rate $\gamma^{*}$ depends on $\sigma$, and how this translates into a bimodal revenue, as in the original monopoly pricing problem. Again, there is a switch from low to high pricing as the variance grows beyond a certain threshold.
Figures LABEL:figmain1hQ and LABEL:figmain1hoptQ show that increasing $h$ or decreasing $\beta$ also forces switching from low to high pricing. Penalizing waiting more heavily indeed makes low pricing with high demand less profitable. And why is it that restricting valuations leads the seller to post high prices and vice versa? Allowing larger valuations increases the set of valuation distributions, providing more opportunity to concentrate mass on low values while meeting the mean constraint. This leaves the seller no other option than selecting a low price. Similarly, preventing the adversary choosing the worst-case distribution from assigning high valuations forces the seller to post a high price. These are universal insights that hold beyond the examples of monopoly pricing and delay-prone services considered in this paper.
@@ 7 Conclusions and outlook
Transitioning from low to high pricing in response to increased market volatility has not been previously documented in the existing literature on robust pricing. When next to the mean, the seller is only aware of the standard deviation, this leads to relatively lower prices that decrease with risk. Similarly, relying exclusively on valuation bounds results in risk-insensitive pricing. The distinctive switch from low to high pricing occurs only when the seller possesses, in addition to the mean and valuation bounds, sufficiently precise information about the standard deviation. This specific combination of factors contributing to the phenomenon might explain why it has been overseen so far. Another reason might be the dominance of low pricing strategies in prior research, coupled with a focus on the intuition that prices should decrease with the level of risk or volatility.
As illustrated in the example of Beta-distributed valuations discussed in the introduction, instances of large or extreme volatility translate into worst-case scenarios, akin to all-or-nothing scenarios. When these scenarios become sufficiently likely, the seller shifts strategies from low pricing and aggressive market penetration to high prices tailored exclusively for those users with the highest valuations. This strategic shift emphasizes the imperative to tailor pricing strategies to accommodate different levels of market uncertainty, particularly in instances of substantial market volatility. We anticipate that the transition to higher pricing due to increased risk, as elucidated in this study via a basic yet parsimonious pricing model, holds broader applicability in more generalized or detailed models and real-world settings. We have demonstrated this already for the setting of delay-prone services.
Exploring the occurrence of low-high pricing in robust settings with partial demand information captured in terms of ambiguity sets is worthwhile to investigate in more detail. Ambiguity sets not explored in this paper include those conditioned on the first three or more moments, or distance-based ambiguity sets centered around a reference distribution, with distance measured in terms of metrics like $\phi$-divergence, Kullback-Leibler distance and Wasserstein distance. While necessitating a distinct mathematical approach, we anticipate that these alternative ambiguity sets may also yield instances of low-high pricing, particularly for sufficiently high levels of ambiguity.
Acknowledgements. This research was supported by NWO Vici grant 202.068. We also thank Alex Suzdaltsev for exchanging ideas about some unpublished results in Suzdaltsev (2018).
Appendix A Proof of Lemma 3.1

### v2

@@ 5 Low and high pricing in service operations management
Monopoly pricing as considered so far considers a firm seeking to maximize the revenue function $p\mathbb{P}(X\geq p)$ where $\mathbb{P}(X\geq p)$ represents the demand function, i.e., the probability that a consumer is willing to pay a price $p$ for some product. When transitioning from products to services, the analysis becomes more intricate. In service systems, the demand function is not fixed but emerges from the decisions of rational customers who evaluate both the price and the expected waiting time. Customers will opt into the service only if the utility they derive exceeds the combined cost of the price and the waiting time. This creates a demand function that is endogenously determined through equilibrium analysis, shaped by the interplay of individual decisions and the operational characteristics of the system. This added complexity makes pricing for services a more challenging problem, as it requires integrating pricing strategies with queueing models to account for congestion and customer behavior. We address robust pricing for services by applying the same max-min framework, focusing on information sets defined by the mean, variance, and range of market characteristics. As will become clear, the core research thread remains consistent: we observe shifts from lower to higher pricing strategies as the market becomes more volatile or heterogeneous. While the resulting price functions and thresholds are less explicit and more involved than in standard monopoly pricing, they underline the universality of the main insight: a robust seller should shift from high to low pricing as the market becomes sufficiently predictable—or conversely, switch to high pricing when the market becomes more uncertain.
5.1 Heterogeneous valuation
Consider a monopolistic service firm that operates as an $M/M/1$ queue. Potential customers arrive according to a Poisson process with arrival rate $\lambda$. Customers who decide to join the queue are served in order of arrival by a single server. Service times are exponentially distributed with rate $\theta$. Following the approach in e.g. Stidham (1992), Cachon and Feldman (2011), Haviv and Randhawa (2014), customers have individual service valuations that are modeled as i.i.d. random variables. In traditional models, the valuation distribution is assumed to be known to the system manager, an assumption that we will relax below. Customers know their valuation but cannot observe the queue length, and decide whether to join or not only based on their assessment of the expected waiting times. Such queues are called unobservable queues and widely studied in the field of rational queueing theory; see Hassin and Haviv (2003) for an extensive overview.
For notational consistency, denote the random valuation of service by $X$, so that a customer only joins the queue when this $X$ exceeds the sum of price $p$ and holding costs $hW(\gamma)$ with $W(\gamma)$ some continuous function that increases with $\gamma$, where a standard choice would be $W(\gamma)=\frac{\gamma}{\theta(\theta-\gamma)}$, the expected waiting time in an $M/M/1$ queue with effective arrival rate $\gamma$ and service rate $\theta$. The parameter $h$ is the cost of waiting per time unit and measures the customer’s delay sensitivity. Let $\mathbb{P}$ denote the distribution of $X$. An arbitrary customer then joins with probability $\mathbb{P}(X>p+hW(\gamma))$. (For mathematical convenience we use a strict inequality >> instead of $\geq$. Note that by Lemma 3.1 this will not affect the worst-case tail bound for the ambiguity sets considered in this paper.) Since customers arrive according to a Poisson process with rate $\lambda$, the effective arrival rate $\gamma=\gamma(p,\mathbb{P})$ solves
@@ 5.1 Heterogeneous valuation
A revenue-maximizing seller seeks to solve $\sup_{p}p\cdot\gamma(p,\mathbb{P})$ and we consider the robust setting where the seller does not know the distribution $\mathbb{P}$ of $X$, but instead only knows that the valuation distribution lies in the ambiguity set $\mathcal{P}$. Taking a pessimistic approach, the seller then sets a deterministic price that maximizes the expected revenue under the worst-case scenario.
It is not guaranteed that (33) has a solution for given $p$ and $\mathbb{P}\in\mathcal{P}$. We therefore set $\gamma(p,\mathbb{P})=\infty$ if no solution to (33) exists. We assume that for every $p$, there is at least one $\mathbb{P}\in\mathcal{P}$ for which $\gamma(p,\mathbb{P})<\infty$. This is in particular true for ambiguity sets satisfying the two conditions in Theorem 5.1, as becomes clear from the proof. We define
@@
and then the objective of a robust seller is to solve $\sup_{p}\,p\cdot\gamma^{*}(p,\mathcal{P})$.
Theorem 5.1 (Robust Endogenous Demand I).
@@
    $\inf_{\mathbb{P}\in\mathcal{P}}\mathbb{P}(X>p+hW(x))$ is attained by some distribution $\mathbb{P}_{x}\in\mathcal{P}$ for every $x$.
@@
The proof can be found in Appendix D.1. Lemma 3.1 shows that Theorem 5.1 applies to all ambiguity sets considered in this paper: The function (6) is always continuous in $p$, and worst-case distributions are identified in the proof of Lemma 3.1 in Appendix A.1. Take as example the ambiguity set $\mathcal{P}(\mu,\beta)$ of all distributions with given mean $\mu$ and maximal valuation $\beta$.
@@
Lemma 3.1 with $\underaccent{\bar}{\sigma}=0$ and $\bar{\sigma}=\sqrt{\mu(\beta-\mu)}$ then yields $\inf_{\mathbb{P}\in\mathcal{P}(\mu,\beta)}\mathbb{P}(X>x)=\frac{\mu-x}{\beta-x}$ for $x\in(0,\mu)$, and hence, by taking $x=p+hW(\gamma^{*})$, we have that $\gamma^{*}$ solves $\frac{\mu-p-hW(\gamma^{*})}{\beta-p-hW(\gamma^{*})}=\frac{\gamma^{*}}{\lambda}$. For $W(\gamma)=\frac{\gamma}{\theta(\theta-\gamma)}$ the worst-case arrival rate $\gamma^{*}$ can be obtained in closed-form. Moreover, for this case, we can also derive an explicit upper bound on $\gamma^{*}$ using the fact that $\mu-p-hW(\gamma^{*})$ needs to be nonnegative. This implies that $\gamma^{*}\leq\gamma_{\max}\mathrel{\mathop{:}}=\theta-\frac{\theta h}{\theta\mu+h-\theta p}$ and this bound becomes exact in the large market size limit $\lambda\to\infty$. This also gives an upper bound $p\gamma_{\max}$ on the worst-case revenue, which is maximized by $p_{\max}\mathrel{\mathop{:}}=\mu-\frac{\sqrt{h^{2}+\theta h\mu}-h}{\theta}$. Pricing in this service system depends on the delay function $W(\cdot)$, the service rate $\theta$, and the additional parameter $h$. Observe that for $h=0$ we retrieve the classical monopoly pricing problem. See Figure LABEL:plotAB for an example. Knowing the mean and maximal value thus leads to a unique optimal price that decreases with $\theta$. As in the original monopoly pricing problem with information on the mean and maximal value, there is no low-high pricing phenomenon. The result, however, is interesting in its own right, as distribution-free pricing for these service systems has not been covered in the literature.
We now return to the main theme of this paper and examine the pricing strategy a seller should adopt when there is information about the variation of the valuation distribution, and thus, the demand uncertainty. Specifically, we assume the seller has knowledge of the mean $\mu$, the support bound $\beta$, and the standard deviation $\sigma$. While the framework extends to cases where the seller only knows that $\sigma$ lies within a certain interval, for clarity of presentation, we focus on the scenario where $\sigma$ is precisely known. Theorem 5.1 implies that the worst-case tail bound, as established in Lemma 3.1, can be utilized to solve the max-min pricing problem given by $\sup_{p}p\gamma^{*}$, assuming $\underaccent{\bar}{\sigma}=\bar{\sigma}$. Figure LABEL:FigmainhQ3 illustrates how the optimal arrival rate $\gamma^{*}$ varies with $\sigma$, leading to a bimodal revenue structure reminiscent of the original monopoly pricing problem. Notably, as standard deviation increases beyond a certain threshold, a transition occurs from low to high pricing, highlighting a fundamental low-to-high pricing phenomenon. Furthermore, we show another switching effect in terms of the support upper bound. As shown in Figures LABEL:figmain1hQ and LABEL:figmain1hoptQ, a decrease in $\beta$ also triggers a switch from low to high pricing. We capture these observations in the following theorem:
Theorem 5.2 (Price Jumps in Queues I).
Assume that $W$ depends on $\sigma$ and $\beta$ solely through $\gamma^{*}$. Consider the optimal price solving $\sup_{p}p\gamma^{*}$ with the distribution of the valuation $X$ contained in $\mathcal{P}(\mu,\sigma,\beta)$. There exist unique thresholds $\sigma^{*}$ and $\beta^{*}$ such that the optimal pricing switches from low pricing for $\sigma\leq\sigma^{*}$ and $\beta\geq\beta^{*}$ to high pricing for $\sigma\geq\sigma^{*}$ and $\beta\leq\beta^{*}$.
Theorem 5.2 is proved in Appendix D.2. Let us elaborate on its implications.
Why does restricting valuations through $\beta$ lead the seller to post higher prices, and vice versa? Allowing larger valuations broadens the set of possible valuation distributions, enabling the adversary to concentrate mass on low valuations while still satisfying the mean constraint. This forces the seller to choose a low price. Conversely, restricting the adversary’s ability to assign high valuations compels the seller to adopt a high-price strategy. These insights are universal and extend beyond the specific settings of monopoly pricing for products and delay-prone services discussed in this paper.
5.2 Heterogeneous delay cost
In rational queueing theory, customer behavior is often modeled based on differences in how much they value the service itself, influencing their willingness to pay for priority or join the queue. This approach was adopted in Section 5.1 and works well in scenarios where customers derive varying benefits from the service. For example, diners at a luxury restaurant may pay more for faster seating to enhance their experience, while concertgoers might purchase VIP tickets for shorter wait times or expedited entry. An alternative perspective focuses on differences in delay costs rather than service valuation. Here, the service is assumed to have equal value for all customers, but individuals vary in their sensitivity to waiting. Those with urgent tasks or high opportunity costs experience greater discomfort from delays, often opting for faster, more expensive options like express lanes or priority services. Meanwhile, customers with lower delay costs are more willing to wait. Examples include airline check-ins, where passengers with tight schedules prioritize speed, or healthcare, where critical patients receive faster treatment due to higher delay costs.
The delay cost rate thus reflects heterogeneity among customers, and accurately characterizing the distribution of delay-cost rates is essential for devising effective pricing strategies.
Much of the existing literature assumes specific forms for the distribution of delay costs. Several works focus on discrete distributions that model multiple types of customers (Ha 2001, Plambeck 2004, Afèche 2013, Ata and Peng 2018). Others investigate continuous distributions, such as uniform delay-cost rates (Gilland and Warsing 2009, Yu et al. 2013). Empirical distributions are derived in Lu et al. (2013) from supermarket sales data and Aksin et al. (2013) use a log-normal distribution based on call center data. Some studies explore broader classes of delay-cost distributions under general regularity conditions (Cao et al. 2019, Nazerzadeh and Randhawa 2018). In contrast to all above works, we adopt a distributionally robust approach, avoiding assumptions about a specific distribution. Instead, we consider all potential distributions within an ambiguity set and explore the delay-cost approach for the robust max-min setting.
Assume that the delay cost rate $H$, measured in dollars per unit time, is a positive random variable. Arriving customers cannot observe queue lengths, but customers do receive an expected delay announcement $W(\gamma)$. Given this information, the customer decides to join by comparing the utility of receiving service $R$ with the cost of waiting. A customer whose delay cost rate is $h$ experiences a disutility of joining of $p+hW(\gamma)$ and joins if $p+hW(\gamma)<R$. The fraction of customers joining will thus be $\mathbb{P}\left(H\leq\frac{R-p}{W(\gamma)}\right)$. Since customers arrive according to a Poisson process with rate $\lambda$, the effective arrival rate $\gamma_{H}=\gamma_{H}(p,\mathbb{P})$ solves $\gamma_{H}=\lambda\mathbb{P}\Big{(}H\leq\frac{R-p}{W(\gamma_{H})}\Big{)}$. Equivalent to (34), we define
$\gamma^{*}_{H}(p,\mathcal{P}):=\inf_{\mathbb{P}\in\mathcal{P}}\gamma_{H}(p,\mathbb{P}),$ (36)
and the seller’s objective $\sup_{p}\,p\cdot\gamma^{*}_{H}(p,\mathcal{P})$. The following theorem then gives a robust bound on the demand:
Theorem 5.3 (Robust Endogenous Demand II).
Let $p>0$ and ambiguity set $\mathcal{P}$ be given, and assume that:
  • (i)
    $g(x)\mathrel{\mathop{:}}=\inf_{\mathbb{P}\in\mathcal{P}}\mathbb{P}(H<\frac{R-p}{W(x)})$ is continuous.
  • (ii)
    $\inf_{\mathbb{P}\in\mathcal{P}}\mathbb{P}(H<\frac{R-p}{W(x)})$ is attained by some distribution $\mathbb{P}_{x}\in\mathcal{P}$ for every $x$.
Then $\gamma^{*}_{H}=\gamma^{*}_{H}(p)$ defined in (36) is the unique solution to the equation
$\gamma^{*}_{H}=\lambda\cdot\inf_{\mathbb{P}\in\mathcal{P}}\mathbb{P}\Big{(}H<\frac{R-p}{W(\gamma^{*}_{H})}\Big{)}.$ (37)
The proof is provided in Appendix D.3. A notable difference between (35) and (37) is the requirement of the supremum rather than the infimum of the tail bound. This follows from the relation
$\inf_{\mathbb{P}\in\mathcal{P}}\mathbb{P}\Big{(}H<\frac{R-p}{W(\gamma^{*}_{H})}\Big{)}=1-\sup_{\mathbb{P}\in\mathcal{P}}\mathbb{P}\Big{(}H\geq\frac{R-p}{W(\gamma^{*}_{H})}\Big{)}.$
We derive this tail bound in Appendix A.2 for $\mathcal{P}=\mathcal{P}(\mu,\underaccent{\bar}{\sigma},\bar{\sigma},\beta)$. We also observe a price jump phenomenon.
Theorem 5.4 (Price Jumps in Queues II).
Assume that $W$ depends on $\sigma$ and $\beta$ solely through $\gamma^{*}_{H}$. Consider the optimal price solving $\sup_{p}p\gamma^{*}_{H}$ with the distribution of the valuation $H$ contained in $\mathcal{P}(\mu,\sigma,\beta)$. There exist unique thresholds $\sigma^{*}$ and $\beta^{*}$ such that the optimal pricing switches from low pricing for $\sigma\leq\sigma^{*}$ and $\beta\geq\beta^{*}$ to high pricing for $\sigma\geq\sigma^{*}$ and $\beta\leq\beta^{*}$.
The proof is provided in Appendix D.4. Observe that the assumption in Theorem 5.4 is satisfied for the expected waiting time in an $M/M/1$ queue.
@@ 6 Conclusions and outlook
The transition from low to high pricing in response to increased market volatility has not been previously documented in the robust pricing literature. When the seller knows only the mean and variance, prices tend to be lower and decrease with risk. Likewise, relying solely on valuation bounds leads to risk-insensitive pricing. The distinctive shift from low to high pricing occurs only when the seller, in addition to knowing the mean and valuation bounds, has sufficiently precise information about the variance. This specific combination of factors may explain why this phenomenon has been overlooked. Another contributing factor could be the dominance of low-pricing strategies in prior research, which has largely focused on the intuition that prices should decrease with risk or volatility.
We have shown that high volatility leads to worst-case, all-or-nothing market scenarios. When these scenarios become sufficiently probable, a robust seller recognizes that the market is strongly segmented. As market uncertainty—and hence potential segmentation—grows, the seller shifts from low pricing and aggressive market penetration to high prices targeting only the highest-valuation users. This strategic shift highlights the necessity of adapting pricing strategies to different levels of market uncertainty, particularly in volatile conditions. We expect that the transition to higher pricing in response to increased risk, as demonstrated in this study through a simple yet insightful pricing model, extends to more general settings. We have explored this phenomenon in delay-prone services, applying our robust max-min analysis to two classical queueing models with rational customers. In these settings, the system operator similarly transitions from low to high pricing when either customer willingness to pay or delay sensitivity becomes too uncertain. Compared to monopoly pricing, pricing services introduces additional challenges, as solving for the worst-case market involves an implicit characterization of the worst-case revenue, which then must be maximized.
The primary goal of this paper was to establish the low-high pricing effect as a consequence of increased segmentation in the worst-case market. To formally prove these results, we adopted a distributionally robust approach, where the seller solves a max-min optimization problem to identify the worst-case market among all distributions with a known mean, maximum valuation, and dispersion (the ambiguity set). Ambiguity sets beyond the scope of this paper include those incorporating the first three or more moments or those defined using distance metrics, such as $\phi$-divergence, Kullback-Leibler distance, or Wasserstein distance. Empirically driven ambiguity sets, which integrate an increasing number of percentiles (based on past pricing responses), also offer an interesting avenue for future research. While requiring a different mathematical framework, we anticipate that these alternative ambiguity sets may also give rise to low-high pricing, particularly under high levels of ambiguity.
For the queueing model, Theorems 5.1 and 5.3 are more general than strictly necessary for the objectives of this paper. While our focus is on the M/M/1 queue with mean-variance ambiguity, these theorems extend to more general queueing models—where the delay function $\gamma\mapsto W(\gamma)$ could, for instance, represent the expected waiting time in an M/G/1 queue—and to broader ambiguity sets. For example, Theorem 5.1 holds for all distributions with a specified mean and support, with the optimal price expressed in closed form. As such, this paper introduces the first robust analysis of rational queues, opening up many directions for future research. A particularly promising avenue is to explore similar models, potentially incorporating heterogeneous valuations and delay sensitivities, under different ambiguity sets.
Appendix A Tail bound results
A.1 Proof of Lemma 3.1